OpenStudy (anonymous):

I tried to solve: X^2 + 24x + 21 = 0 by completing the square method, but when i need to take square root i get a -ve no under radical, but when i solve using a quadratic formula i get a answer.

OpenStudy (anonymous):

x^2+24x+21=(x+12)^2-144+21=(x+12)^2-123 (x+12)^2-123=0 => x+12=sqrt(123)=>x=sqrt(123)-12 or x+12=-sqrt(123)=>x=-sqrt(123)-12

OpenStudy (anonymous):

Thank you very much i was not subtracting the -144 after adding it.

OpenStudy (anonymous):

no problem